OpenStudy (anonymous):

Express your answer in lowest terms. -2/9*(-3/14)=

OpenStudy (anonymous):

giving medals

OpenStudy (amorfide):

\[\frac{ a }{ b} \times \frac{ c }{ d }= \frac{ a times c }{ b times d }\]

OpenStudy (amorfide):

that should read \[\frac{ a \times c }{ b \times d }\]

OpenStudy (anonymous):

ok now what

OpenStudy (amorfide):

well let a/b = -2/9 and c/d= -3/14

OpenStudy (amorfide):

just multiply the top numbers, multiply the bottom numbers

OpenStudy (anonymous):

so i just multiply ok thx

OpenStudy (anonymous):

-2 -3 -- --- = 6 / 126 or 1/21 9 14

OpenStudy (anonymous):

ok thanks